# Closed mean curvature flows with asymptotically conical singularities

### v1

@@ Abstract.
In this paper, we prove that for any asymptotically conical self-shrinker, there exists a closed hypersurface such that the mean curvature flow starting from it develops a singularity modeled on the given shrinker. The main technique is the Ważewski box argument, used by Stolarski [Sto] in the proof of the corresponding theorem in the Ricci flow case.
@@ 1. Introduction
We consider a family of immersions $\mathbf{x}\colon M^{n}\times I\to\mathbb{R}^{n+k}$ such that $\partial_{t}\mathbf{x}=\mathbf{H}$ where $\mathbf{H}$ is the mean curvature vector of the immersed submanifold $M_{t}=\mathbf{x}(M,t).$ This family of immersions is called a mean curvature flow. Understanding the singularity formations in a mean curvature flow is a fundamental question. More precisely, given a mean curvature flow $M_{t}$ developing a spacetime singularity at $(\textbf{0},0)\in\mathbb{R}^{n+k}\times\mathbb{R},$ we can consider a sequence of rescaled mean curvature flows $M^{i}_{t}:=\lambda_{i}M_{\lambda_{i}^{-2}t}$ using any sequence $\lambda_{i}\to\infty.$ Based on Huisken’s monotonicity [H90] and Brakke’s compactness [B78], White [W94] and Ilmanen [I95] proved that $M_{t}^{i}$ weakly converges to a self-similar limit flow, called a tangent flow, after replacing $\lambda_{i}$’s with a subsequence. When the limiting flow is smooth, it is generated by a self-shrinker, a submanifold satisfying $\textbf{H}=-\frac{x^{\perp}}{2}.$ Uniqueness of tangent flows, a fundamental issue for singularity analysis, has been discussed for certain shrinkers [S14, CM15, CM19, Z20, CS, LSS, LZ, CSSZ]. For related references on uniqueness of tangent cones of minimal submanifolds, one can see [AA, S83, H97, S08, CM14].
@@
Given an asymptotically conical shrinker $\Sigma^{n}\subset\mathbb{R}^{n+1}$, there exists a closed embedded hypersurface $M\subset\mathbb{R}^{n+1}$ such that the mean curvature flow $M_{t}$ starting from $M$ develops a type I singularity at time $1$ at the origin modeled on $\Sigma$.
This confirms the belief in the existence of closed mean curvature flows having asymptotically conical tangent flows. It then has some applications based on earlier works. For example, Sun-Xue [SX2] proved that given a closed mean curvature flow with an asymptotically conical singularity, after a generic initial perturbation, the perturbed mean curvature flow avoids the asymptotically conical singularity. Their theorem suggests that it is hard to see an asymptotically conical singularity in closed mean curvature flows. In fact, it is expected and has been proven in the surface case (cf. [CCMS, CCS, SX1, SX2, BK23]) that all non-cylindrical singularities are not generic.
@@
This paper is organized as follows. In Section 2, we provide preliminary notations and results for asymptotically conical self-shrinkers and talk about the evolution equations of related immersions. In Section 3, we give an explicit construction of an embedded doubling of an asymptotically conical shrinker. In Sections 4 and 5, we talk about the interior estimates and constructions of barrier functions to get the preservation and the improvement of $C^{2}$ bounds. In Section 6, we use the Ważewski box argument to finish the proof.
@@ Acknowledgement
The authors are grateful to Prof. Toby Colding and Prof. Bill Minicozzi for their constant support and several inspiring conversations. They would also like to thank Max Stolarski for helpful conversations. During the project, Lee was partially supported by NSF Grant DMS 2005345 and Zhao was supported by NSF Grant DMS 1812142 and NSF Grant DMS 1811267.
@@ 2.1. Stability operators on asymptotically conical shrinkers
(2.1) $C=\lim\limits_{t\to 0}\sqrt{-t}\Sigma$
@@
(2.2) $\left|\nabla^{i}A\right|\leq\frac{C_{j}}{|x|^{j+1}}$
@@
The corresponding $L^{2}$ and $H^{k}$ spaces will be denoted by $L^{2}_{W}(\Sigma)$ and $H^{k}_{W}(\Sigma).$ As in the Ricci flow case ([Sto, Lemma 2.16]), the inclusion map given by $H^{1}_{W}(\Sigma)\to L^{2}_{W}(\Sigma)$ is a compact embedding. This is rigorously proven in [BW17, Appendix B] (cf. [CS, LZ]). In [CS], it is also shown that $L\colon H^{2}_{W}\left(\Sigma\right)\to L^{2}_{W}(\Sigma)$ is a bounded operator. Hence, elliptic theories implies that the spectrum of $L$ behaves as nicely as we expected.
@@
The next thing we need is the nice decaying property of these eigenfunctions. Similar to the eigenfunctions of the weighted Lichnerowicz operator, one expects that the $L$-eigenfunctions also decay in a suitable manner. This part follows from a general result proven by Colding-Minicozzi in [CM21]. A simpler case for $\mathcal{L}$-eigenfunctions was dealt with by Colding-Minicozzi in [CM20].
@@
Now we apply [CM21, Theorem 3.9] by considering $f=\frac{|x|^{2}}{4},$ $b=|x|=r,$ and $\psi=Ch^{2}$ in their work and obtain the desired estimate for $h.$ Note that the $\lambda$-eigenspace is finite-dimensional, so we can find a constant dependeing on $\lambda$ working for all $\lambda$-eigenfunctions. ∎
@@ 3. Constructions of doubling of asymptotically conical shrinkers
As in (2.1), denoting the asymptotic cone of $\Sigma$ by $C$, from [W14, CS], we know that outside a large compact set, we can parameterize $\Sigma\setminus B_{R}(0)$ as
@@
over $C$, where $\theta$ is the parametrization for the link of $C$ and n is the normal vector of $C$ at $(r,\theta)$.
@@
The first step is to move the link to a hemisphere. As the link of $C$ is a codimension one submanifold of $S^{n-1}$, we can pick a vector $\mathbf{v}\in S^{n-1}$ such that $-\mathbf{v}$ is not contained in the link of $C$. Initially, we have $F_{1}(r,\theta)=r\theta$ for $r>2R.$ We take
@@ 5. Barrier functions
This proposition seems very technical, but it enables us to easily control the $C^{0}$ norm of h using this barrier function. By combining Proposition 4, Lemma 4, Lemma 4, and Lemma 4, we can prove the global $C^{2}$-improvement theorem following the same argument in [Sto]. Note that here we need $C^{1}$-bounds instead of $C^{0}$-bounds in the mean curvature flow setting.
@@ 6. Ważewski’s box argument
and thus, from the definition of the box, ww obtain

### v2

@@ Abstract.
In this paper, we prove that for any asymptotically conical self-shrinker, there exists an embedded closed hypersurface such that the mean curvature flow starting from it develops a singularity modeled on the given shrinker. The main technique is the Ważewski box argument, used by Stolarski in the proof of the corresponding theorem in the Ricci flow case. As a corollary, our construction, combined with the works of Angenent–Ilmanen–Velázquez and Chodosh–Daniels-Holgate–Schulze, implies the existence of fattening level set flows starting from smooth embedded closed hypersurfaces. This addresses a question posed by Evans–Spruck and De Giorgi.
@@ 1. Introduction
We consider a family of immersions $\mathbf{x}\colon M^{n}\times I\to\mathbb{R}^{n+k}$ such that $\partial_{t}\mathbf{x}=\mathbf{H}$ where $\mathbf{H}$ is the mean curvature vector of the immersed submanifold $M_{t}=\mathbf{x}(M,t).$ This family of immersions is called a mean curvature flow. Understanding the singularity formations in a mean curvature flow is a fundamental question. More precisely, given a mean curvature flow $M_{t}$ developing a spacetime singularity at $(\textbf{0},0)\in\mathbb{R}^{n+k}\times\mathbb{R},$ we can consider a sequence of rescaled mean curvature flows $M^{i}_{t}:=\lambda_{i}M_{\lambda_{i}^{-2}t}$ using any sequence $\lambda_{i}\to\infty.$ Based on Huisken’s monotonicity [H90] and Brakke’s compactness [B78], White [W94] and Ilmanen [I95] proved that $M_{t}^{i}$ weakly converges to a self-similar limit flow, called a tangent flow, after replacing $\lambda_{i}$’s with a subsequence. When the limit flow is smooth, it is generated by a self-shrinker, a submanifold satisfying $\textbf{H}=-\frac{x^{\perp}}{2}.$ Uniqueness of tangent flows, a fundamental issue for singularity analysis, has been discussed for certain shrinkers [S14, CM15, CM19, Z20, CS, LSS, S23, LZ, CSSZ]. For related references on uniqueness of tangent cones of minimal submanifolds, one can see [AA, S83, H97, S08, CM14].
@@
Given an asymptotically conical shrinker $\Sigma^{n}\subset\mathbb{R}^{n+1}$, there exists an embedded closed hypersurface $M\subset\mathbb{R}^{n+1}$ such that the mean curvature flow $M_{t}$ starting from $M$ develops a type I singularity at time $1$ at the origin modeled on $\Sigma$.
This confirms the belief in the existence of closed mean curvature flows having asymptotically conical tangent flows. This finding has applications, drawing on previous research. For instance, Chodosh–Choi–Mantoulidis–Schulze [CCMS] proved that for a closed mean curvature flow with an asymptotically conical singularity, a generic initial perturbation ensures the perturbed flow avoids the asymptotically conical singularity. Their theorem indicates that encountering an asymptotically conical singularity in closed mean curvature flows is unlikely. Indeed, it is anticipated that all non-cylindrical singularities are not generic, a conjecture confirmed for surfaces [CCMS, CCS, BK23]. Another method addressing this issue, based on local dynamics, was explored by Sun–Xue [SX1, SX2].
Another immediate consequence of Theorem 1 is noteworthy. In [ES], Evans–Spruck conjectured that any level set flow starting from a smooth hypersurface does not fatten (cf. the final sentence in [ES]). Based on the construction by Angenent–Ilmanen–Velázquez [AIV] and the fattening criterion proven by Chodosh–Daniels-Holgate–Schulze [CDS], Theorem 1 implies the following corollary. It provides a negative answer to the question raised by Evans–Spruck (cf. [De, I94, AIC]) by showing that fattening does not necessarily happen instantaneously.
Corollary 1.1.
There exists a fattening level set flow starting from an $n$-dimensional smooth embedded closed hypersurface for $n\in\left\{3,4,5,6\right\}.$ In particular, the flow remains smooth for a short time before fattening occurs.
In [W02], White announced a result joint with Ilmanen about the existence of a compact smooth embedded surface in $\mathbb{R}^{3}$ for which the uniqueness of enhanced varifold solutions of mean curvature evolution fails. He described the proof ideas in [W02, Section 4], which are completely different from our approach to Corollary 1.1. Combining their result with Corollary 1.1, we now have $n$-dimensional embedded closed examples for $n\in\left\{2,3,4,5,6\right\}.$
@@
This paper is organized as follows. In Section 2, we provide preliminary notations and results for asymptotically conical self-shrinkers and talk about the evolution equations of related immersions. In Section 3, we give an explicit construction of an embedded doubling of an asymptotically conical shrinker. In Sections 4 and 5, we talk about the interior estimates and constructions of barrier functions to get the preservation and the improvement of $C^{2}$ bounds. In Section 6, we use the Ważewski box argument to finish the proof of Theorem 1 and prove Corollary 1.1.
@@ Acknowledgement
The authors are grateful to Prof. Toby Colding and Prof. Bill Minicozzi for their constant support and several inspiring conversations. They would like to thank Letian Chen, Otis Chodosh, Alec Payne, Max Stolarski, and Ao Sun for valuable discussions, and to Alec Payne for bringing the work of Angenent, Ilmanen, and Velázquez to their attention and suggesting the application regarding fattening. During the project, Lee was partially supported by NSF Grant DMS 2005345 and Zhao was supported by NSF Grant DMS 1812142 and NSF Grant DMS 1811267.
@@ 2.1. Stability operators on asymptotically conical shrinkers
(2.1) $\mathcal{C}=\lim\limits_{t\to 0}\sqrt{-t}\Sigma$
@@
(2.2) $\left|\nabla^{j}A\right|\leq\frac{C_{j}}{|x|^{j+1}}$
@@
The corresponding $L^{2}$ and $H^{k}$ spaces will be denoted by $L^{2}_{W}(\Sigma)$ and $H^{k}_{W}(\Sigma).$ As in the Ricci flow case ([Sto, Lemma 2.16]), the inclusion map given by $H^{1}_{W}(\Sigma)\to L^{2}_{W}(\Sigma)$ is a compact embedding. This is rigorously proven in [BW17, Appendix B] (cf. [CS, LZ]). In [CS], it is also shown that $L\colon H^{2}_{W}\left(\Sigma\right)\to L^{2}_{W}(\Sigma)$ is a bounded operator. Hence, elliptic theories imply that the spectrum of $L$ behaves as nicely as we expect.
@@
The next thing we need is a nice decaying property of these eigenfunctions. Similar to the eigenfunctions of the weighted Lichnerowicz operator, one expects that the $L$-eigenfunctions also decay in a suitable manner. This part follows from a general result proven by Colding–Minicozzi in [CM21]. A simpler case for $\mathcal{L}$-eigenfunctions was dealt with by Colding–Minicozzi in [CM20].
@@
Now we apply [CM21, Theorem 3.9] by considering $f=\frac{|x|^{2}}{4},$ $b=|x|=r,$ and $\psi=Ch^{2}$ in their work and obtain the desired estimate for $h.$ Note that the $\lambda$-eigenspace is finite-dimensional, so we can find a constant depending on $\lambda$ working for all $\lambda$-eigenfunctions. ∎
@@ 3. Constructions of doubling of asymptotically conical shrinkers
As in (2.1), denoting the asymptotic cone of $\Sigma$ by $\mathcal{C}$, from [W14, CS], we know that outside a large compact set, we can parameterize $\Sigma\setminus B_{R}$ by
@@
over $\mathcal{C}\setminus B_{R}$, where $\theta$ is the parametrization for the link of $C$ and n is the normal vector of $\mathcal{C}$ at $(r,\theta)$.
@@
The first step is to move the link to a hemisphere. As the link of $\mathcal{C}$ is a codimension one submanifold of $S^{n-1}$, we can pick a vector $\mathbf{v}\in S^{n-1}$ such that $-\mathbf{v}$ is not contained in the link of $\mathcal{C}$. Initially, we have $F_{1}(r,\theta)=r\theta$ for $r>2R.$ We take
@@ 5. Barrier functions
This proposition seems very technical, but it enables us to easily control the $C^{0}$ norm of h using this barrier function. By combining Proposition 4, Lemma 4, and Lemma 4, we can prove the global $C^{2}$-improvement theorem following the same argument in [Sto]. Note that here we need $C^{1}$-bounds instead of $C^{0}$-bounds in the mean curvature flow setting.
@@ 6. Ważewski’s box argument
and thus, from the definition of the box, we obtain
@@
Using Theorem 1 and the results from [AIV, CDS], we can prove Corollary 1.1.
Proof of Corollary 1.1.
Angenent–Ilmanen–Velázquez constructed $n$-dimensional asymptotically conical shrinkers such that level set flows starting from their asymptotic cones fatten for $n\in\left\{3,4,5,6\right\}$ in [AIV, Theorem 1]. [^1] The proof of Theorem 1 in [AIV] was provided in Section 3 (for proving Theorem 5) and Section 5.1 there. Take such a shrinker $\Sigma$ and let $\mathcal{C}$ be its asymptotic cone. Theorem 1 implies that there exists a smooth closed hypersurface $M\subseteq\mathbb{R}^{n+1}$ such that the level set flow $M_{t}$ starting from $M$ develops a singularity at $({\bf 0},0)\in\mathbb{R}^{n+1}\times\mathbb{R}$ whose tangent flow is modeled on $\Sigma.$ Thus, $M_{0}$ is a hypersurface with a singularity at the origin modeled on the cone $C$ (cf. [CS, Corollary 1.2]). Since the level set flow starting from $\mathcal{C}$ fattens, the main result in [CDS] implies that the level set flow $M_{t}$ fattens for positive time. ∎